OpenStudy (anonymous):

Is (a^4+b^4) the same as (a+b)^4

hartnn (hartnn):

no.

OpenStudy (anonymous):

\[(a+b)^4=a^4+4a^3b^2+6ab+4a^2b^3+b^4\] we see a^4+b^4 on the right so they are not the same

OpenStudy (anonymous):

for eg take a=2 and b=3 from (i)=a^4+b^4=2^2+3^4=16+81=97 while from (ii)=(a+b)^4=5^4=625 :)

OpenStudy (anonymous):

do you know the FOIL method? if so: is (a^2 + b^2) the same as (a+b)^2 no because the 2nd is (a+b)*(a+b) = a^2+2ab+b^2 same principle